OpenStudy (anonymous):

(4(y+2)^(1/2)-2y(y+2)^(-1/2))/(y+2) simplify

OpenStudy (anonymous):

This is what you're working with, right? \[\frac{4(y+2)^{1/2}-2y(y+2)^{-1/2}}{y+2}\]

OpenStudy (campbell_st):

split the problem \[\frac{4(y + 2)^{\frac{1}{2}}}{(y + 2)^1}\] and \[-\frac{ 2y(y + 2)^{-\frac{1}{2}}}{(y + 2)^1}\] just subtract the powers

OpenStudy (campbell_st):

oops since you have terms in (y +2) just subtract the powers

OpenStudy (anonymous):

OK, yeah, that'll work. Carry on.

OpenStudy (anonymous):

what do you mean by "just subtract the powers"?

OpenStudy (campbell_st):

when dividing you can subtract powers for the same base a^3/a^4 = a^1